OpenStudy (anonymous):

Solve the following inequality. 1/3x^3/4(x − 4)^−2/3 + 3/4x^−1/4(x − 4)^1/3 < 0

OpenStudy (anonymous):

I might have wrote it alittle confusing let me try to clear it up

OpenStudy (squirrels):

Thanks c:<

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }x ^{3/4}(x-4)^{-2/3}+\frac{ 3 }{ 4 }x ^{-1/4}(x-4)^{1/3}<0\]

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

\(\large \frac{ 1 }{ 3 }x ^{3/4}(x-4)^{-2/3}+\frac{ 3 }{ 4 }x ^{-1/4}(x-4)^{1/3}<0\) \(\large \frac{ 1 }{ 3 }x ^{3/4}(x-4)^{-2/3}<-\frac{ 3 }{ 4 }x ^{-1/4}(x-4)^{1/3}\) \(\large \frac{ 1 }{ 3 }(x ^{\frac{-9}{8}}(x-4)^{\frac{-2}{3}})<-\frac{ 3 }{ 4 }(x ^{\frac{3}{4}}(x-4)^{\frac{1}{3}} )\)

OpenStudy (ikram002p):

sry last step i made a typo

OpenStudy (ikram002p):

\(\large \frac{ 1 }{ 3 }(x ^{\frac{-9}{8}}(x-4))^{\frac{-2}{3}}<-\frac{ 3 }{ 4 }(x ^{\frac{3}{4}}(x-4))^{\frac{1}{3}} \)

OpenStudy (jdoe0001):

\(\bf \large \cfrac{1}{3}x^{\frac{3}{4}}\cdot (x-4)^{-\frac{2}{3}}+\cfrac{3}{4}x^{-\frac{1}{4}}\cdot (x-4)^{\frac{1}{3}} < 0 \quad ?\)

OpenStudy (ikram002p):

\(\large(x ^{\frac{-9}{8}}(x-4)^{\frac{-2}{3}})<-\frac{ 9 }{ 4 }(x ^{\frac{3}{4}}(x-4)^{\frac{1}{3}} )\) raise both side to power 3 \(\large (x ^{\frac{-9}{8}}(x-4)^{\Huge{-}2})<-\frac{ 27 }{ 64 }(x ^{\frac{3}{4}}(x-4)^{1} )\)

OpenStudy (ikram002p):

lol are u getting this ?

OpenStudy (ikram002p):

@jdoe0001 it can be solved :D

OpenStudy (anonymous):

okay so i figured out how to do it and that is not correct. you have to leave everything to the right of the less than sign and leave zero by itself. then put the negative numbers under the denominator and find a common factor. thank you for trying though!

OpenStudy (anonymous):

negative exponents* i should say